answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the expected outcome of melodic intonation therapy?

The expected outcome after completion of the MIT sequence is increased communication through production of intelligible word groups. Patients are typically able to form short sentences of 3-5 words, but more complex communication may also be possible

You randomly draw 1 card from a standard deck of 52 and 1 card from another standard deck Whats the probability that the sum of the 2 cards is 7?

If you consider an Ace as 1, then first make sure that the first card is less than a 7. There are 6 numbers below 7, and 24 cards below 7. The second card has to be an exact number, which is always 4/52. Both cards together make a 6/13 * 1/13 = 6/169 probability.

What is the difference between probability and statistics?

Sl.no Probability Statistics 01 Probability deals with predicting the likelihood of future events statisticsinvolves the analysis of the frequency of past events 02 Probability is primarily a theoretical branch of mathematics, which studies the consequences of mathematical definitions Statistics is primarily an applied branch of mathematics, which tries to make sense of observations in the real world. 03 probability theory enables us to find the consequences of a given ideal world Statistical theory enables us to measure the extent to which our world is ideal.

Where can you buy jelly marbles?

The jelly marbles are made and sold by Steve Spangler Science. They are $10 for about 1000 marbles (50 g) and kit.

See related link "Jelly Marbles" for more information

Is the cyclic redundancy check better than hamming code?

It depends on what you are doing.

The cyclic redundancy check will only detect an error, while the hamming code can also correct many types of errors. However to perform this correction the extra error detection parity bits required in hamming code are many more than the bits needed for cyclic redundancy check, per data byte being checked.

Normally cyclic redundancy check is done on large block of data that can be resent or retried to get the correct block of data (e.g. telecommunication channels, disk sectors).

Normally hamming code is done on individual bytes or words of computer memory.

What is the meaning of standard yield?

Start with good output 54,000 kg

the standard yield 80%

Calculate the standard input

54000/.80 = 67,500 kilograms

What is the value of rolled uncirculated rolled 2000 P Sacajawea dollars?

More than 767 MILLION coins were struck in the first year of issue at the Philadelphia Mint. Original bags and rolls are so common most dealers don't buy them. But you may find someone that wants them.

What is the probability of choosing a queen or a king from a standard deck of 52 playing cards?

If one card is picked at random from a well shuffled deck, the answer is 8/52 = 2/13

What is the probability of a card higher than a five clubs appearing in a suit of clubs?

The probability of a card higher than a five in any particular suit, when that suit is the only suit in the deck of 13, assuming the Ace is high, is 9 in 13, or about 0.6923. If the Ace is low, then the probability is 8 in 13, or about 0.6154.

If you draw from a full deck of 52 cards, and stipulate that you have to draw a club, then the probabilities become 9 in 52, or about 0.1731 (Ace high) or 8 in 52, or about 0.1538 (Ace low).

Can an event E have a probability of -1?

No; probability is between 0 and 1 inclusive.

How many different 5 letter arrangements can be formed from the letters in the word Danny?

The number of 5 letter arrangements of the letters in the word DANNY is the same as the number of permutations of 5 things taken 5 at a time, which is 120. However, since the letter N is repeated once, the number of distinct permutations is one half of that, or 60.

When data is normally distributed which test can be taken?

The answer depends on what is being tested: the t-test, F-test, Chi-square, Z-test are all commonly used with the Normal distribution. There are many others.

Trending Questions
What is the P TTT on three flips of a coin? A child rolls a 6-sided die 6 times What is the probability of the child rolling exactly two fours? What is the meaning of ''bucket-and-spade'' holiday in tenby? Box A has 3 red marbles and 2 yellow marble Box B has 2 red marble and 1 yellow If you have to pick a red marble to win and can't look in the box which box would give you the best chance? What is the probability that you do not pick a picture card? How many different three-letter initials are there that begin with an A? How many times must you flip a coin before there is at least a 50 percent probability that you will get at least 3 heads? How likely are you to pull a red or yellow marble out of a bag of 3 green 8 red 8 yellow and 3 black marbles? 1The odds in favor of A solving mathematical problem are 3 to 4 and the odds against B solving the problems are 5 to 7 Find the probability that the problem will be solved by at least one of them? What about accidental sampling? What is the theorical probability of rolling a 4 after rolling a number cube 72 times? What is the probability of rolling a number greater than ten when 2 6 sided dice are thrown? How many possible hands are there for getting 4 cards in the same suit with the other cards in any other random suit? What was the first pr? What is the event if the probability will equal 1? How many different arrangements of the letters in the word COLOR are there? How do you determining desired outcomes? Can a z-score be converted into a probability? What is the probability of 2 drownings in a city of 200000? How many possible outcomes when tossing 3 coins?